### v1

# Cross-modal Cognitive Consensus guided Audio-Visual Segmentation

Abstract
Audio-Visual Segmentation (AVS) aims to extract the sounding object from a video frame, which is represented by a pixel-wise segmentation mask. The pioneering work conducts this task through dense feature-level audio-visual interaction, which ignores the dimension gap between different modalities. More specifically, the audio clip could only provide a Global semantic label in each sequence, but the video frame covers multiple semantic objects across different Local regions. In this paper, we propose a Cross-modal Cognitive Consensus guided Network (C3N) to align the audio-visual semantics from the global dimension and progressively inject them into the local regions via an attention mechanism. Firstly, a Cross-modal Cognitive Consensus Inference Module (C3IM) is developed to extract a unified-modal label by integrating audio/visual classification confidence and similarities of modality-specific label embeddings. Then, we feed the unified-modal label back to the visual backbone as the explicit semantic-level guidance via a Cognitive Consensus guided Attention Module (CCAM), which highlights the local features corresponding to the interested object. Extensive experiments on the Single Sound Source Segmentation (S4) setting and Multiple Sound Source Segmentation (MS3) setting of the AVSBench dataset demonstrate the effectiveness of the proposed method, which achieves state-of-the-art performance.
Index Terms:
Audio-visual segmentation, Cross-modal cognitive consensus, Semantic-level consistencyI Introduction
Interesting object segmentation is fundamental for high-efficiency multimedia analysis. In recent years, object segmentation has been well explored for various visual signals, which strive to extract all objects or stuff from different granularity including the semantic segmentation [1, 2, 3], instance segmentation [4, 5, 6], and panoptic segmentation [7, 8]. Meanwhile, these great efforts also bring us sweet trouble. That is, are all the objects/stuff interesting or necessary for the users in analyzing the prevalent multimedia data, which typically contains both audio and visual modalities?
In [9], Zhou et al. made a new exploration toward Audio-Visual Segmentation (AVS), which focuses on extracting the sounding objects from a video frame based on their pixel-wise correspondence. This new task provides multiple modalities to facilitate more specific interesting object extraction and a dense feature-level audio-visual interaction framework is proposed to achieve this target. Despite the success in integrating multi-modal features [9], it is still challenging to well establish the semantic correspondence between audio and visual modalities due to the dimension gap. More specifically, the audio clip could only provide a Global semantic label in each sequence, but the video frame covers multiple semantic objects across different Local regions.
To address the aforementioned issues, we propose a novel Cross-modal Cognitive Consensus guided Network (C3N). The schematic diagram of our method is shown in Fig 1. The C3N aligns audio and video from the global dimension via semantic-level information (i.e. classification confidence and semantic similarities) and integrates the cross-modal cognitive consensus with local visual regions through an attention mechanism. Firstly, the proposed Cross-modal Cognitive Consensus Inference Module (C3IM) feeds the extracted visual and audio features into independent heads to obtain classification confidence and convert the modality-specific labels into embedding vectors to calculate the cross-modal semantic similarities. We multiply audio/visual confidence scores and semantic similarities to infer the cross-modal cognitive consensus and extract a unified-modal label. Next, feed the unified-modal label back to hierarchical layers of the visual backbone as the global semantic-level guidance for highlighting local features corresponding to the object of interest in a channel-spatial attention manner through the parameter-free Cognitive Consensus guided Attention Module (CCAM). Finally, input the refined features into a cross-modal feature fusion module and adopt a segmentation head to integrate multi-modal features for the final segmentation.
The major contributions are concluded as follows:
We design a novel Cross-modal Cognitive Consensus Inference Module (C3IM) leveraging semantic-level audio/visual confidence and similarities of labels to infer cross-modal cognitive consensus and extract a unified-modal label for the subsequent segmentation guidance.
The parameter-free Cognitive Consensus guided Attention Module (CCAM) is proposed to highlight the visual local feature elements corresponding to the sounding object with the guidance of the global unified-modal label in a channel-spatial attention manner.
We propose a Cross-modal Cognitive Consensus guided Network (C3N) for the AVS task, whose key components are C3IM and CCAM. The results of the experiments show that C3N outperforms other methods and achieves state-of-the-art performance on the AVSBench dataset.
II Related Work
II-A Visual Object Segmentation
Researchers have proposed many methods for segmenting images or videos that only rely on visual input. Semantic Segmentation requires pixel-level category assignment of the given image, for which various CNN-based [1, 10, 11] and Transformer-based [12, 13, 14] models are well-designed with high accuracy. Moreover, Instance Segmentation [4, 15, 16, 17] aims to distinguish all objects into individual instances. Panoptic Segmentation [7, 18, 8] unifies the above two tasks, i.e. discriminating all objects and stuff in an image, and assigns semantic labels.
Video Object Segmentation (VOS) focuses on segmenting an object in a sequence of video frames. Semi-supervised VOS (SVOS) [19, 20, 21, 22, 23, 24] means the annotation in the first frame is given as guidance during inference. OSVOS [19] relies on a pre-trained fully convolutional network. MaskTrack [25] and RGMP [20] are propagation-based methods, which make predictions based on the mask of the previous frame. VideoMatch [26] and A-GAME [27] perform pixel-level match of frames to be segmented with the reference frame. FEELVOS [21] and CFBI [22] adopt the embedding learning method and perform pixel-level and instance-level matching between frames. SSTVOS [24] leverages the sparse attention mechanism to extract temporal-spatial relevance. Unsupervised VOS (UVOS) [28, 29, 30, 31, 32] means no human guidance information is available during model inference. Wang et al. [30] and COSNet [31] use attention-based methods to emphasize the correlations between video frames. Hu et al. [29], Li et al. [33], and Ren et al. [32] introduce additional motion cues during segmentation.
Despite the remarkable achievement of the visual object segmentation, the absence of information from other modalities prevents it from emphasizing objects of interest efficiently.
II-B Sound Source Localization
Recently, researchers attempt to incorporate audio to extract the visual object of interest such as the Sound Source Localization (SSL) task aims to locate the sounding object. SSL evolves from Audio-Visual Correspondence (AVC), whose pioneering work [34] fuses audio and visual features to measure audio-visual consistency. With the growth of deep learning, many SSL methods [35, 36, 37, 38, 39, 40, 41, 42] have been proposed. AVOL-Net [37] takes the middle layer output of the AVC network as the localization. Senocak et al. [36] adopts contrastive learning for audio-visual feature-level alignment. Chen et al. [40] and Lin et al. [43] improve localization performance through hard negative mining. Some recent methods [38, 39, 41, 42, 44] try to localize in multi-source scenes. Hu et al. [38, 44] propose clustering-based methods to obtain category-aware cluster centers and partition multiple objects. Qian et al. [39] tackle this task by using the class-activated map (CAM). Song et al. [41] propose a negative-free localization method achieved by visual data augmentation to mine explicit positive examples. Afouras et al. [45], Cheng et al. [46] and Tian et al. [47] propose general audio-visual frameworks, which are suitable for various downstream tasks including SSL. However, SSL generates heat maps for coarsely localizing sounding objects, which lack fine-grained descriptions.
Zhou et al. [9] propose Audio-Visual Segmentation (AVS) task that predicts pixel-level masks for sounding objects, and an AVS baseline utilizes feature-level interactions for multi-modal fusion. However, due to the lack of semantic-level guidance, it remains hard to align audio and video explicitly.
II-C Label Embedding
Word embedding means mapping distinct words into representational vectors to allow the neural networks to learn the contents and semantic relationship of words, which can be divided into static methods such as word2vec [48], Glove [49], and contextural-based methods such as ELMo [50], GPT [51] and BERT [52]. These powerful methods help researchers incorporate text semantics to facilitate the reasoning process.
In many cases, researchers convert label texts into embedding vectors and combine them with extracted features to enrich the framework’s semantic-level knowledge [53, 54, 55, 56, 57, 58, 59]. Zhang et al. [54] point out the importance of label information and design multiple label embedding-based models. LEAM [55] and EXAM [57] introduce interaction mechanisms to extract input-label relevance clues. GILE [56] proposes a generalized input-label strategy to strengthen the model’s performance on unseen classes. HARNN [60] incorporates recurrent layers with attention mechanism and models multi-layers dependencies. Xiong et al. [58], Cai et al. [61], and Wang et al. [62] introduce label embedding into BERT [52] or its variants to further enhance the semantic-level sensibility of the model. LTTA-LE [63] is a truncation-based method that leverages the label embedding to filter the redundant information of the long text and reduce the text length. In addition, the effectiveness of label embedding is also demonstrated in visual tasks [64, 65] and zero-shot tasks [53, 66].
Label embedding is currently adopted mainly on the above single-modal tasks such as image classification and text classification, while exploration in global semantic-level audio-visual alignment based on label embedding is still limited.
III Proposed Method
The overall architecture of our Cross-modal Cognitive Consensus guided Network (C3N) is illustrated in Fig. 2. The network takes an audio clip and video frames as inputs, where , and are the length, height, and width of the input video frame sequence. Firstly, we use the audio encoder and visual backbone to extract audio features and hierarchical visual features , where is the feature dimension and . Then, a Cross-modal Cognitive Consensus Inference Module (C3IM) uses pre-trained heads to compute the audio/visual classification confidence and calculate the similarities between audio and visual labels for extracting the semantic-level unified-modal label. Next, the Cognitive Consensus guided Attention Module (CCAM) highlights the local feature elements of the sounding object with the guidance of the unified-modal label. Finally, feed the refined visual features into the cross-modal feature fusion module and make predictions through the segmentation head.
III-A Cross-modal Cognitive Consensus Inference
We aim to explicitly extract global semantic-level alignment information to improve the performance of the AVS framework. Therefore, we propose a novel Cross-modal Cognitive Consensus Inference Module (C3IM) to exploit semantic-level audio/visual classification confidence and labels to infer the audio-visual cognitive consensus.
The visual backbone is initialized with weights pre-trained on ImageNet [67], a large-scale image classification dataset totaling 1000 classes. And the audio encoder is pre-trained on AudioSet [68] with 527 audio classes. Such a diversity of classes applies to the semantic description of audio/visual objects in most scenarios. In practice, we freeze the parameters of the audio and visual classification heads. Firstly, as shown in Fig. 3, for every-second audio feature and the highest level visual feature , we compute the corresponding classification confidence scores:
where and are pre-trained weights of the classification heads. and are classification confidence scores, where and are confidence of the audio class and visual class.
Then, to bridge the semantic gap between distinct modalities, we measure the similarities of audio and visual labels and construct a semantic similarity matrix. The audio labels and visual labels are denoted as and respectively. We convert the audio and visual labels to a unified label embedding space by means of SpaCy, an advanced library for natural language processing to calculate similarities between label words or phrases. The audio and visual label embeddings are represented as and respectively, where the embedding dimension is 300. The similarity between the audio and visual labels is defined by the following equation:
where denotes the Frobenius norm, and , are the row and column indexes of the semantic similarity matrix .
Finally, based on the computed audio and visual classification confidence scores , and the similarity matrix , a confidence re-weighted matrix can be calculated by the following equation:
where and are balance coefficients, which are set to 0.1. Values in the re-weighted matrix can also consider as cross-modal cognitive consensus degrees. With the inferred cognitive consensus, semantically relevant audio-visual objects can be identified to guide the following segmentation.
III-B Cognitive Consensus guided Attention
After the cross-modal cognitive consensus inference, we feed the semantically consistent unified-modal label to the hierarchical layers of the visual backbone. In addition, we propose a parameter-free Cognitive Consensus guided Attention Module (CCAM), which injects class-activated weights into visual features to highlight the local feature elements corresponding to the sounding object.
We search for the maximum value of and obtain its row and column indexes and . Specifically, and refer to the index of the audio and visual classes with the highest cognitive consensus score. Inspired by the ideology of Grad-CAM [69], we calculate the loss for the visual class and perform the backpropagation to compute the corresponding gradients. Different from Grad-CAM, which computes class-activated heat maps, we aim to obtain the weights corresponding to the unified-modal label with the highest cognitive consensus. This process can be defined as:
where denotes classification logit of the visual class, denotes feature map activations, denotes the index of the channel. In practice, we take the last layer of each stage as the activation layer and the class-activated weights of stage is denoted as .
To integrate the class-activated weights with the initial visual features for highlighting the local feature elements corresponding to high cognitive consensus, we propose a parameter-free CCAM, whose structure is shown in the bottom right of Fig 2. Without any learnable parameters, CCAM performs channel-spatial attention to the initial visual features and output refined features with the guidance of inferred global semantic-level unified-modal label, which can be expressed as:
where denotes sigmoid activation function, denotes element-wise multiplication with boardcast mechanism. By means of the channel-spatial attention mechanism, CCAM incorporates semantic-level cognitive consensus weights with feature-level visual representations to guide the model focus on the semantically consistent visual object.
III-C Cross-modal Feature Fusion
For cognitive consensus refined visual features , Atrous Spatial Pyramid Pooling (ASPP) [10] modules are used to help the perception at multiple receptive fields. We denote the ASPP-processed features as .
After getting the visual feature , feed it with the audio feature into a cross-modal Non-local block [70, 9], which utilizes 3-D convolutional layers to model the feature-level temporal-spatial dependencies and captures the inter-modal correlations. We first map and repeat into . Then, the cross-modal feature-level interaction process at stage can be formulated as follows:
where , , and denote 3-D convolutional layers. is the number of pixels of the feature map, which is treated as a normalization constant. is the cross-modal attention matrix, which measures the pixel-wise correlations between audio and visual features. denotes the multi-modal feature of the stage.
III-D Segmentation Head
We combine the hierarchical multi-modal features in a simple top-down manner by progressively upsampling the higher-level feature and integrating it with the lower-level feature rich in detail information. This process can be formulated as follow:
Then the final prediction masks can be obtained by feeding into a fully convolutional head.
To optimize the parameters of the proposed model. We adopt the binary cross entropy (BCE) loss as the loss function:
where denotes the ground truth masks.
IV Experiments
IV-A Experimental Settings
IV-A1 Dataset
The proposed method is evaluated on the AVSbench [9] dataset for audio-visual segmentation. To the best of our knowledge, there is only one open-source dataset (AVSBench) in the AVS field. Videos from AVSBench are collected from YouTube and each video is cropped to 5 seconds. Video frames at the end of each second are extracted. Depending on the number of sounding objects in the video, the AVSBench dataset is divided into a Single-Source subset and a Multi-Source subset. The Single-Source subset contains 4,392 videos from 23 categories. Note that only the first sampled frame is annotated for videos in the train split, while all sampled frames of the val/test split are annotated, totaling 10,852 annotations. The Multi-Sources subset contains 424 videos and 2,120 annotated pixel-level segmentation masks. All sampled frames of the train/val/test split have annotation since the sounding objects may change over time. As with the pioneering work [9], the following experiments are carried out under the semi-supervised Single Sound Source Segmentation (S4) setting and the fully supervised Multiple Sound Source Segmentation (MS3) setting.
IV-A2 Implementation Details
We implement the proposed method using PyTorch [71]. We use Swin Transformer (Swin) [72] and Pyramid Vision Transformer v2 (PVTv2) [73] as the visual backbones. The channel sizes of the four stages are for Swin and for PVTv2. The visual backbones, including the corresponding classification heads, are pre-trained on ImageNet-1K [67] dataset and the weights of the classification heads are frozen. For audio encoders, the widely adopted VGGish [74] and PANNs [75] pre-trained on AudioSet [68] are used to extract audio features. Since the VGGish does not have a classification head, we trained an additional head for classification using audio features from AudioSet. All of the video frames are resized to without any data augmentation or post-process methods, and the audios are clipped to 1-second splits. The channel size of is set to 256. In all experiments, the models are optimized by Adam [76] and the learning rate is 0.0001. The batch size is set to 4 and the number of training epochs is 20 for the semi-supervised S4 setting and 60 for the fully supervised MS3 setting. We train our model on an RTX 3090 GPU.
IV-A3 Evaluate Metrics
We adopt the mean intersection-over-union (mIoU) and F-score as evaluation metrics to measure the performance of the proposed method quantitatively. The mIoU is calculated by dividing the intersection area by the union area of the predictions and ground truths and taking the average value of the whole dataset and F-score 111, is set to 0.3, which remains the same as in [9]. considers both precision and recall.
IV-B Comparison with State-of-the-Art Methods
Method | Backbone | mIoU () | F-score () | ||
S4 | MS3 | S4 | MS3 | ||
LVS[40] | ResNet-18 2 | 37.9 | 29.5 | 51.0 | 33.0 |
MSSL[39] | ResNet-18+CRNN | 44.9 | 26.1 | 66.3 | 36.3 |
3DC[77] | ResNet-152 (3D) | 57.1 | 36.9 | 75.9 | 50.3 |
SST[24] | ResNet-101 | 66.3 | 42.6 | 80.1 | 57.2 |
iGAN[78] | Swin | 61.6 | 42.9 | 77.8 | 54.4 |
LGVT[79] | Swin | 74.9 | 40.7 | 87.3 | 59.3 |
Baseline[9] | ResNet-50+VGGish | 72.8 | 47.9 | 84.8 | 57.8 |
Baseline[9] | PVTv2+VGGish | 78.7 | 54.0 | 87.9 | 64.5 |
Baseline-Re | ResNet-50+VGGish | 72.7 | 46.6 | 83.2 | 56.3 |
Baseline-Re | PVTv2+VGGish | 78.8 | 54.0 | 87.8 | 65.1 |
C3N (Ours) | Swin+VGGish | 77.5 | 53.5 | 86.9 | 64.7 |
C3N (Ours) | Swin+PANNs | 77.8 | 54.4 | 87.0 | 65.0 |
C3N (Ours) | PVTv2+VGGish | 80.0 | 56.9 | 88.8 | 67.4 |
C3N (Ours) | PVTv2+PANNs | 80.4 | 57.0 | 88.9 | 68.0 |
We use the mIoU and F-score metrics to quantitively compare our method with state-of-the-art methods [40, 39, 77, 24, 78, 79, 9] under the S4 and MS3 settings. We show the performance of the AVS baseline and the authors’ re-implementation of other SSL/VOS/Salient Object Detection (SOD) methods on AVSBench reported in [9]. In addition, we run the authors’ open-source code on our experimental devices under the default settings in [9] to obtain the quantitive performance of the baseline models. The same code running conditions are also employed to get the results of our C3N.
Table I shows the results of the test set. The proposed C3N with PVTv2 [73] as the visual backbone and VGGish [74] or PANNs [75] as the audio encoder achieves remarkable performance and outperforms the baseline method by a large margin. For the simple S4 setup, compared with the baseline with PVTv2 as the visual backbone, the gains achieved by C3N (PVTv2+VGGish) and C3N (PVTv2+PANNs) are 1.2, 1.6 in mIoU and 1.0, 1.1 in F-score, respectively. For the difficult MS3 setting with multiple sounding objects, thanks to the guidance of semantic-level cognitive consensus, C3N (PVTv2+VGGish) and C3N (PVTv2+PANNs) achieve higher performances with a large margin of 2.9, 3.0 mIoU, and 2.3, 2.9 under F-score metric.
We also adopt Swin Transformer [72], a universal visual backbone, to evaluate the general effectiveness of our cognitive consensus-based method. Our C3N with Swin Transformer and VGGish as backbones achieves 77.5 mIoU and 86.9 F-score under the S4 setting and 53.5 mIoU and 64.7 F-score under the MS3 setting. While the results of C3N (Swin+PANNs) are slightly higher. The performance of C3N with Swin as the backbone is close to or even surpasses the PVTv2 version baseline under the MS3 setting, and far higher than the ResNet version under the S4 and MS3 settings.
IV-C Ablation Study
IV-C1 Analysis of key components
Swin+VGGish | ||||||
AF | CC-V | CC | mIoU () | F-score () | ||
S4 | MS3 | S4 | MS3 | |||
75.1 | 47.1 | 85.2 | 56.4 | |||
✓ | 75.6 | 47.1 | 85.5 | 57.6 | ||
✓ | 76.1 | 48.7 | 86.2 | 58.9 | ||
✓ | 76.4 | 50.2 | 85.7 | 61.0 | ||
✓ | ✓ | 77.5 | 53.5 | 86.9 | 64.7 | |
Swin+PANNs | ||||||
AF | CC-V | CC | mIoU () | F-score () | ||
S4 | MS3 | S4 | MS3 | |||
75.4 | 48.3 | 85.5 | 58.0 | |||
✓ | 75.9 | 49.0 | 85.9 | 58.9 | ||
✓ | 76.4 | 49.1 | 85.8 | 59.8 | ||
✓ | 76.7 | 51.6 | 86.1 | 61.4 | ||
✓ | ✓ | 77.8 | 54.4 | 87.0 | 65.0 | |
PVTv2+VGGish | ||||||
AF | CC-V | CC | mIoU () | F-score () | ||
S4 | MS3 | S4 | MS3 | |||
77.8 | 51.0 | 87.1 | 62.5 | |||
✓ | 78.5 | 51.1 | 87.4 | 61.0 | ||
✓ | 78.7 | 52.5 | 87.6 | 63.4 | ||
✓ | 79.2 | 54.3 | 87.9 | 64.9 | ||
✓ | ✓ | 80.0 | 56.9 | 88.8 | 67.4 | |
PVTv2+PANNs | ||||||
AF | CC-V | CC | mIoU () | F-score () | ||
S4 | MS3 | S4 | MS3 | |||
77.7 | 49.8 | 86.8 | 61.2 | |||
✓ | 78.2 | 50.9 | 87.2 | 60.9 | ||
✓ | 79.0 | 51.7 | 88.0 | 62.4 | ||
✓ | 79.2 | 54.3 | 87.9 | 64.9 | ||
✓ | ✓ | 80.4 | 57.0 | 88.9 | 68.0 |
We conduct ablation study experiments with all four backbone combinations. In the experiments, we investigate the effects of audio-visual feature-level interaction, cognitive consensus relies merely on visual modality, and the full audio-visual cognitive consensus. The results are shown in Table II.
The first row shows the performance without audio features and cognitive consensus, i.e. with only video frames as input. Under all four backbone combinations, as in [9], visual-only input does not cause a performance drop by a large margin for the S4 setting while causing a distinct performance drop under the MS3 setting. It indicates that for videos with multiple sounding sources, the introduction of the audio signal and alignment between audio-visual modalities are especially essential to the final prediction. Moreover, in the second row, we introduce the ”CC-V”, which means that the labels with the highest visual semantic confidence are fed into the framework. The results show that only relying on semantic-level guidance from a single modality cannot bring noticeable improvements.
Then we add semantic-level cognitive consensus in the third row. It means the audio-visual semantic-level cognitive consensus is introduced but does not yet include audio features. The audio-visual semantic-level cognitive consensus leads to performance gains for all four groups. In detail, under the MS3 setting, mIoU increases by 1.6, 0.8, 1.5, and 1.9 respectively, and F-score increases between 0.9 to 2.5. For the S4 setting, mIoU increases by around 1 under each backbone combination, and F-score also increases by various points. Furthermore, compared with the results in the fourth row that add audio features, the results in the third row are competitive under the S4 setting. The above results demonstrate that the model performance can be improved through the cross-modal cognitive consensus-based method.
We incorporate audio features and conduct feature-level interaction in the fourth row. The results represent that introducing audio features to the visual input is effective. The reason is that powerful pre-trained audio encoders can capture the dense information of the audio signal. Moreover, the non-local blocks extract cross-modal pixel-wise correlations, which is beneficial to the integration of the two modalities. Nevertheless, such cross-modal interaction is inexplicit due to the absence of semantic-level cognitive consensus, which leads to errors during the segmentation.
Finally, we add both audio features and semantic-level cognitive consensus to the model, i.e. the proposed C3N. It can be noticed that with the cognitive consensus-based cross-modal alignment, the performance of the model improves remarkably under all four backbone settings. In particular, under the MS3 setting, the gains of mIoU are 3.3, 2.8, 2.6, 2.7 respectively and F-score increases by 3.7, 3.6, 2.5 and 3.1. For the S4 setting, mIoU and F-score metrics generally rise by around 1 in all four experimental groups. The above results demonstrate the audio-visual cognitive consensus inference and feedback in our method facilitates explicit cross-modal alignment and improves AVS model performance.
Swin+PANNs | |||||
mIoU () | F-score () | ||||
S4 | MS3 | S4 | MS3 | ||
0.10 | 0.10 | 76.8 | 58.8 | 86.1 | 70.2 |
0.25 | 0.25 | 76.7 | 58.8 | 86.1 | 70.2 |
0.10 | 0.25 | 76.7 | 58.7 | 86.3 | 69.8 |
0.25 | 0.10 | 76.4 | 58.8 | 85.8 | 69.7 |
IV-C2 Analysis of Hyperparameters
The two hyperparameters involved in the C3N framework are balance coefficients and in Equation (4). Thue, we conduct multiple settings for the two parameters on the validation set of the S4 and MS3 scenarios under the Swin+PANNs backbone setting. The detailed quantitive results are shown in Table III.
Experimental results show that the mIoU and F-score metrics under different parameter settings (i.e. ==0.10, ==0.25, =0.10 and =0.25, =0.25 and =0.10) fluctuate within 0.5 on the validation set, which is a relatively small range. Finally, we choose ==0.10 as the balance coefficients to construct the confidence re-weighted matrix.
IV-D In-depth Analysis
IV-D1 Qualitative analysis
In Fig. 4, we present two AVS examples on the test set of the AVSBench dataset for a qualitative comparison between the baseline method and our C3N. We show segmentation masks of two baseline models (i.e. ResNet-50+VGGish version and PVTv2+VGGish version) and two C3N models (i.e. Swin+VGGish version and PVTv2+VGGish version) with the same audio encoder to ensure objective comparison.
As shown in the left panel, two guitars are simultaneously played. The Res+VGGish version and PVTv2+VGGish version baselines fail to accurately depict both sounding guitars and predict many false positive areas. However, our proposed method uses CCAM to refine the visual features with the guidance of semantic-level cognitive consensus. As a result, our C3N (Swin+VGGish) and C3N (PVTv2+VGGish) segment the sounding object’s contour more precisely.
For the right sample of Fig. 4, the video shows a child playing with a dog as the child continuously makes giggling sounds and the dog is silent. The example is hard because the color of the dog is significantly different from the background, which may mislead the model into incorrectly segmenting the dog as well. The baseline method (Res+VGGish) even completely ignores the sounding child sometimes. The baseline (PVTv2+VGGish) and C3N (Swin+VGGish) segment the dog and child simultaneously. The reason why our C3N (Swin+VGGish) makes false segmentation of the dog is that the Swin generates unbalanced confidence scores and overly attends to the dog in the image, despite the dog and kid should be equally treated without audio guidance, which causes the unified-modal label shifts to the dog. However, the C3N (PVTv2+VGGish) overcomes distraction from the other object and accurately segments the sounding child with tolerable false positive areas.
IV-D2 Effect of cross-modal cognitive consensus inference
The C3IM integrates the classification confidence of distinct modalities with similarities between modality-specific labels to extract the semantic-level unified-modal label. In Fig. 5, we present two visualization examples of the cross-modal cognitive consensus inference process. For ease of illustration, we only show top-5 confidence classes for audio and visual modalities. Note that the audio and visual confidence scores correspond to and in Equation. (4). The highest semantic similarity value, confidence re-weighted value, and the corresponding unified-modal label are framed.
In the top example, the image shows a man playing the guitar, and the audio is the sound of the guitar. Column (b) shows that the visual classification head can not tell if it is an electric guitar or an acoustic guitar, and the man’s clothing also has high confidence. In column (c), the audio-visual label similarities are calculated to establish initial semantic-level correlations. In column (d), the confidence re-weighted matrix is obtained by multiplying audio/visual confidence and cross-modal label similarities. It is noticeable that the visual class corresponding to the maximum value is acoustic guitar, which is consistent with the input audio and visual information. Based on the unified-modal label, the model perceives the object of interest. After that, columns (e) and (f) show the model accurately segments the sounding acoustic guitar.
In the bottom example, the sounding object is the violin. This example is more difficult because other prominent targets are present in the visual scene, such as humans and the piano. The sounding violin occupies only a small part of the frame. Column (b) presents the grand piano has the highest confidence and the model incorrectly regard the violin as an oboe. However, after combining it with the audio confidence and semantic similarities, the inferred cognitive consensus still emphasizes the correct violin, fiddle class, which is shown in column (d). According to such guidance information, the model correctly segments the violin. The above visualization results confirm the validity of our C3IM.
IV-D3 Effect of cognitive consensus guided Attention
The Cognitive Consensus guided Attention Module (CCAM) aims to highlight the local feature elements of the sounding object based on inferred cross-modal cognitive consensus. For comparison, we visualize the attention matrices within the audio-visual attention module of the AVS baseline [9] and the spatial attention map within the CCAM of the proposed C3N. In practice, we remove the non-linear activations of CCAM.
In Fig. 6, the heat maps of attention maps of the AVS baseline and the CCAM are shown in the third and fourth rows, respectively. Note that in the fourth row, only the semantic-level alignment of audio and visual modalities is performed and audio features are absent, while the third row shows the audio-visual feature-level attention. In the left example, the violin produces sounds and other objects keep silent. In the third row, although the AVS baseline locates the violin, other irrelevant regions are also highlighted to varying degrees. Nevertheless, in the fourth row, only the sounding violin is emphasized in the spatial attention map of the CCAM. In the right sample, the audio-visual situation is more complex, as two instrument sounds are mixed, and the scene switches in the video. It requires the model to accurately perceive the sounding objects in distinctive scenes. The AVS baseline incorrectly highlights drummers, however, in the fourth row, with the help of cross-modal cognitive consensus, the CCAM heat maps locate the sounding drum and guitar in the first three frames and the last two frames respectively, which is consistent with ground truths.
The above results demonstrate the CCAM allows accurately perceiving and localizing the sounding object only relying on semantic-level cognitive consensus and facilitates the subsequent cross-modal feature fusion and segmentation.
V Conclusion
In this paper, we have proposed Cross-modal Cognitive Consensus guided Network (C3N), a novel framework for Audio-Visual Segmentation that exploits semantic-level information for explicit cross-modal alignment. We obtain the unified-modal label by integrating audio/visual classification confidence and semantic similarities via a Cross-modal Cognitive Consensus Inference Module (C3IM). In addition, we develop a Cognitive Consensus guided Attention Module (CCAM) to highlight the local features corresponding to the object of interest depending on the global cognitive consensus guidance. Extensive experiments verify the effectiveness of our method and its superiority over state-of-the-art methods.